A football field is 40 yards wide and 120 yards long. Find the distance between opposite corners of the football field. Show your work and explain your reasoning.
Answer by richwmiller(17219)

40^2+120^2=c^2
1600+14400=160000
126.49
hypotenuse of right triangle with sides 40 and 120